OpenStudy (anonymous):

can someone please help me asap 2a^2-71-15/3a^2-8a-3 times 9a^2-1/4a^2-9/a^2+3a-10/2a^2-91+9

OpenStudy (aum):

Without parenthesis it is difficult to say what is numerator and denominator. Can you post either a screenshot of the question or use the equation editor to rewrite the expression so it is clearer?

OpenStudy (anonymous):

\[\frac{ 3x^2 }{ x^2 }- \frac{ x^3 }{ x^3-1 }-\frac{ 2x }{ x-1 }\]

OpenStudy (anonymous):

first find the LCD..